The ratio of two complimentary angles is 7:8. What is the measure of the larger angle?
Found 2 solutions by andecover, stanbon:
Answer by andecover(2)
Answer by stanbon(75887)

The ratio of two complimentary angles is 7:8. What is the measure of the larger angle?
-----
7:8 is the same as 7x:8x where 8x is the measure of the larger angle.
----
Equation:
7x+8x = 90 degrees
15x = 90 degree
x = 6
-----
Larger angle = 8x = 8*6 = 48 degrees
===========
Cheers,
Stan H.
=============